Consider the following sample data:


x 11 7 5 5 4
y 3 10 13 6 11


Click here for the Excel Data File
a.

Calculate the covariance between the variables. (Negative value should be indicated by a minus sign. Round your intermediate calculations to 4 decimal places and final answer to 2 decimal places.)


  Covariance   


b.

Calculate the correlation coefficient. (Round your intermediate calculations to 4 decimal places and final answer to 2 decimal places.)


  Correlation coefficient   
Answer #1

Solution-:

We prepare the following table:

x y x^2 y^2 x*y
11 3 121 9 33
7 10 49 100 70
5 13 25 169 65
5 6 25 36 30
4 11 16 121 44
Total 32 43 236 435 242

  From this table we get

n=5,\sum X=32,\sum Y=43,\sum X^2=236,\sum Y^2=435,\sum XY=242

\bar{X}=\frac{\sum X}{n}=\frac{32}{5}=6.4000 and \bar{Y}=\frac{\sum Y}{n}=\frac{43}{5}=8.6000

Var(X)=\frac{\sum X^2}{n}-\bar{X}^2=\frac{236}{5}-6.4^2=6.2400

Var(Y)=\frac{\sum Y^2}{n}-\bar{Y}^2=\frac{435}{5}-8.6^2=13.0400

SD(X)=\sqrt{Var(X)}=\sqrt{6.24}=2.4980

SD(Y)=\sqrt{Var(Y)}=\sqrt{13.04}=3.6111

(a) Cov(X,Y)=\frac{\sum XY}{n}-\bar{X}*\bar{Y}=\frac{242}{5}-6.4*8.6=-6.64

(b)Corr(X,Y)=r=\frac{-6.64}{(2.4980*3.6111)}=-0.74

There is negative correlation between X and Y